# Privacy Reasoning in Ambiguous Contexts

### v2

@@ 1 Introduction
Personal AI agents are expected to interact with a user to complete tasks on their behalf. To effectively complete a broad variety of tasks in a personalized manner, such agents require access to significant amounts of personal and potentially sensitive user data. This implies that an agent must be able to safeguard this personal information while achieving high utility. To be truly useful, the agent must therefore be able to judge based on the context of the interaction when it is appropriate (as defined in the contextual integrity framework nissenbaum2004privacy; nissenbaum2009privacy) and consistent with a specific user’s expectations to share personal information in interactions with third-party tools and agents.
Prior research mireshghallah2023confaide; salt-nlp2024privacylens has investigated the ability of large language models (LLMs) to maintain the secrecy of sensitive information, with initial studies reporting limited success. More recent findings with advanced LLMs, however, indicate substantial improvements on certain datasets sun2024trustllm. Nevertheless, these existing benchmarks focus on whether an LLM avoids disclosing sensitive information and do not equally address the crucial aspect of utility; that is, whether the model can contextually determine when sharing sensitive information is appropriate.
@@
First, we introduce a comprehensive dataset encompassing both positive (appropriate to share) and negative (inappropriate to share) examples, expanding upon existing datasets that predominantly feature negative instances. Our study then reveals an oversensitivity to prompt variations in LLM decisions—a signal that can indicate complexity in contextual integrity analysis shvartzshnaider2024position. We also find that classification errors frequently stem from underspecified context, leading the models to make implicit assumptions. In instances where model predictions diverge, the model’s internal reasoning is consistent with privacy reasoning we expect to see. However, it leads to a different outcome due to the implicit application of different, yet seemingly plausible, contextual interpretations. We term these scenarios “ambiguous contexts”: situations lacking the contextual details necessary to definitively determine the appropriateness of data sharing.
@@ 2 Background and Related Work
Reasoning and understanding. Nuanced privacy judgments benefit from an understanding of the mental states of other parties, through theory of mind. Such capabilities of LLMs have been studied both generally kim2023fantom; sclar2023minding and in relation to privacy mireshghallah2023confaide. Privacy assessments also rely on complex reasoning, which has been studied extensively through benchmarks in other domains. StrategyQA geva2021aristotle and MuSR sprague2023musr test implicit and multi-step reasoning. Logical reasoning errors have been explored in ZebraLogic yuchen2025zebralogic, the cause of reasoning errors has been investigated in REVEAL jacovi2024chain, and recently the ability to ask clarifying questions in underspecified contexts has been studied in QuestBench li2024questbench on logical puzzles, planning, and math problems. Our work complements benchmarks developed for general reasoning and increases the trustworthiness of these general-purpose models, through measurements of how advancements in general reasoning capabilities translate to improving privacy reasoning and assessments. It also extends the analysis of underspecified context to privacy reasoning.
@@
LLM privacy assessments. Prior work has studied privacy issues related to LLMs from multiple angles. One line of work aims to assess carlini2021extracting; staab2023beyond and mitigate abadi2016deep; li2021large; yu2021differentially leakage of training data. As the capability of LLMs to consume larger contexts has improved lewis2020retrieval, allowing them to perform agent tasks on user data bagdasarian2024airgapagent; ngong2025protecting; zharmagambetov2025agentdam, the research area of inference-time privacy has emerged. Inference-time privacy assessments can generally be categorized as either studies of privacy leakage or privacy awareness huang2025trustworthiness. Privacy leakage measurements, e.g. through red teaming, focus on capturing the refusal to answer wang2023decodingtrust; zhang2024multitrust; sun2024trustllm, and do not assess utility in terms of sharing sensitive data in appropriate contexts. In contrast, privacy awareness assessments are based on probing-style prompts to test the capability of LLMs in various privacy settings, where the appropriateness of sharing data is context-dependent. Since the goal of our benchmark is to capture the impact of context underspecification on privacy assessments, we focus on privacy awareness.
@@
Privacy awareness and contextual integrity. Due to the contextual nature of appropriateness, many privacy awareness assessments of LLMs have been grounded in contextual integrity theory nissenbaum2004privacy; nissenbaum2009privacy, which defines privacy as adherence to context-specific informational norms. Several studies focus on the alignment of models with norms and user preferences li2024goldcoin; li2024privacychecklist; shvartzshnaider2024llm. Multiple such studies focus on building benchmarks for various tasks. The work by Ghalebikesabi et al. ghalebikesabi2024operationalizing develops a benchmark for form-filling. PrivaCI-Bench li2024privacibench focuses on benchmarks for legal compliance. CI-Bench cheng2024ci builds a benchmark for assistant tasks. A concurrent study by Lan et al. lan2025contextual focuses on improving privacy reasoning through reinforcement learning. These papers focus on judging the capability of LLMs in assessing information-sharing appropriateness by also measuring utility, unlike our study that focuses on evaluating the impact of underspecified context on privacy reasoning. Moreover, these benchmarks are either not available publicly, or are focused on specific domains such as law. The two publicly available benchmarks that we use in our work are PrivacyLens salt-nlp2024privacylens and ConfAIde mireshghallah2023confaide, which we describe in detail in Appendix D and expand upon in Section 4.
@@ 3 Problem Statement
We consider an agent, depicted in Figure 1, modeled by an LLM $\phi$, tasked with making a binary judgment about sharing a piece of information (e.g., a medical condition or an appointment time). Such information is served to the model through data structures called scenarios $\mathcal{S_{C}}$, parameterized by a set of contextual fields $\mathcal{C}$, often inspired by the contextual integrity theory nissenbaum2004privacy; nissenbaum2009privacy. The task of the model $\phi$ is then to output $\phi(\mathcal{S_{C}})\in\{\texttt{yes},\texttt{no}\}$, where yes (respectively no) corresponds to the conclusion that it is appropriate (respectively not) to share the data in the scenario $\mathcal{S_{C}}$. See Figure 2 for an example.
@@ 4 Context Ambiguity Analysis
To better evaluate both the privacy and utility of LLMs in probing-style tasks, we use two existing privacy benchmarks: PrivacyLens salt-nlp2024privacylens and ConfAIde mireshghallah2023confaide. PrivacyLens contains 493 examples characterized by five contextual integrity fields: data type, data subject, data sender, data recipient and transmission principle. In ConfAIde, examples describe scenarios where an aware agent tries to disclose a sensitive personal detail about a subject agent to an oblivious agent for a reveal reason. Each of the 270 ConfAIde examples contains seven fields (the subject agent’s aware-agent-relationship and oblivious-agent-relationship, in addition to the five above), which are expanded into short descriptive stories on which privacy judgments are performed. Because the two datasets contain only inappropriate-to-share scenarios (denoted as $\mathcal{S^{-}}$) to measure privacy leakage, we augment them by incorporating instances where data is appropriate to share ($\mathcal{S^{+}}$). This allows us to perform comprehensive assessments of utility. For each $\mathcal{S^{-}}$, we generate a corresponding positive scenario $\mathcal{S^{+}}$ by modifying a subset of its fields. For PrivacyLens, we source these appropriate-to-share fields directly from the dataset, while for ConfAIde we generate them using an LLM. We validate the positive examples manually. The resulting augmented datasets, termed PrivacyLens+ and ConfAIde+, incorporate $\mathcal{S^{+}}$ alongside the original $\mathcal{S^{-}}$, and contain 986 and 540 examples respectively. A detailed description of the dataset generation procedure is found in Appendix D.
@@
Following existing benchmarks mireshghallah2023confaide, we start with a neutral prompt that asks “is it okay” to share the relevant data without providing more information or context. Next, we create a restrictive prompt variant by adding “Please answer the question considering privacy norms in the response” to the neutral prompt, akin to prior work mireshghallah2023confaide. Finally, we create a permissive prompt variant by adding “Please answer the question sharing as much as possible while respecting privacy.” The goal of these variants is to identify how performance changes across distinct input instructions, rather than to develop the best prompt variants for each setting. In addition to these variants that change privacy directives in the prompts, we also evaluate the model’s sensitivity to paraphrasing-style perturbations in Appendix B.2.4. The complete prompts are listed in Appendix E. For all experiments, we set temperature to 0 and the maximum output length to one token (the yes/no judgment).
@@
Table 1 highlights that model performance is highly sensitive to the prompt variant used across datasets. The neutral prompt generally yields high precision (85.6–92.2% for PrivacyLens+ and 98.6–100.0% for ConfAIde+) but lower recall (62.7–75.7% for PrivacyLens+ and 85.6–95.6% for ConfAIde+), indicating models prioritize privacy over utility, achieving high precision at the cost of low recall. The high precision also aligns with prior work finding that models more easily distinguish negative examples sun2024trustllm.
@@ 5 Camber: A Context Disambiguation Framework
The first strategy serves as a baseline and adds arbitrary information to one of the data fields (e.g., clarifying the data subject of a request from “alumnus” to “an alumnus who had been recognized for contributions in environmental science”), helping us measure to what extent any additional privacy-agnostic information is useful. To ensure that the added information remains in agreement with the existing appropriateness label of the example, we prompt the model to generate only neutral information, in line with prior work salt-nlp2024privacylens. In the second baseline strategy, we aim to synthetically add information that is more privacy-related and indicative of appropriateness to further resolve ambiguity (e.g., retrieving information that the student record was marked as “do not disclose”). To ensure that this added information remains in agreement with the appropriateness label, we include the label in the data generation prompt. We call the first strategy label-independent and the second label-dependent. Finally, our third strategy mimics an agent that reasons about which fields in the information request are ambiguous and seeks evidence to reduce this ambiguity (e.g., reasoning that the sender is not authorized to share alumni information externally). We call this the reasoning-guided expansion strategy (Section 5.2). Comparing results from the first two strategies reveals how important it is for agents to clarify privacy-specific aspects compared to asking general questions about the context. The reasoning-guided strategy further intends to shed light on what types of privacy relevant information would be most effective to acquire. To evaluate these strategies and their impact on improving privacy judgments, in Section 5.1 we describe the mechanisms enabling Camber to generalize across both the ConfAIde+ and PrivacyLens+ datasets. A detailed description of the Camber framework can be found in Appendix C.
@@ 5.1 Enabling context disambiguation across datasets
To investigate to what extent context that clarifies the appropriateness label is effective at aiding LLM judgment, we define a second baseline that adds contextual details to $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples according to the definition of the fields in each dataset. These label-dependent contexts are either inappropriate or appropriate, denoted as $\mathcal{N^{-}_{D}}$ and $\mathcal{N^{+}_{D}}$ respectively. Incorporating these contexts yields examples denoted as $\mathcal{S^{-}_{N^{-}_{D}}}$ and $\mathcal{S^{+}_{N^{+}_{D}}}$. For both datasets, we first ask the LLM to generate $\mathcal{N^{-}_{D}}$ (respectively, $\mathcal{N^{+}_{D}}$) for a specific field in $\mathcal{S^{-}}$ ($\mathcal{S^{+}}$) to make the example more inappropriate (respectively, appropriate). For ConfAIde+, we then additionally ask the LLM to integrate the generated context into the original example. It is important to ensure that the expansion does not result in reasoning shortcuts, wherein the added context directly reveals the appropriateness label. We achieve this by prompting the context-generating LLM to avoid directly expanding on the appropriateness label, and using heuristics that remove certain words indicative of the label, in line with prior work salt-nlp2024privacylens. We repeat this procedure for all fields in each dataset, resulting in an additional 2,465 pairs of $\mathcal{S^{-}_{N^{-}_{D}}}$ and $\mathcal{S^{+}_{N^{+}_{D}}}$ examples for PrivacyLens+, and 1,890 pairs for ConfAIde+. The detailed generation procedure can be found in Appendix D.1.2 and D.2.3.
@@ 5.2 Reasoning-guided expansion
Building on extensive research into LLM reasoning geva2021aristotle; sprague2023musr; yuchen2025zebralogic; jacovi2024chain and factors influencing their outputs, we seek to understand appropriateness decisions of LLMs by prompting them to provide concise reasoning alongside each yes/no decision (our prompts can be found in Appendix E). Although elicited reasoning may not perfectly represent internal processes for models, it offers valuable insights into their decision-making chen2025reasoning. This reasoning elicitation step yielded performance comparable to that of models not providing explicit reasoning (P=83.5%, R=74.4%, and $F_{1}$=78.7% for PrivacyLens+, and P=97.6%, R=75.9%, and $F_{1}$=85.4% for ConfAIde+).
To understand model reasoning and identify implicit assumptions, two authors analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; FPs were excluded due to insufficient samples). They examined the prompts and reasoning for those examples—blind to ground-truth labels and LLM judgments—to identify assumptions and develop the nine codes shown in Table 2. Subsequently, using this established codebook, the same two authors independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreement was resolved through discussion to achieve consensus cascio2019team. Table 2 presents the resulting code occurrences across datasets.
@@
To test if codes are more effective at reducing implicit reasoning assumptions, we prompt the LLM to generate label-dependent contexts based on each code (denoted as $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$) for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples respectively in both datasets. Our instruction specifically asks the LLM to first identify the most suitable field to expand based on the code and its definition, and then to expand the selected field following this code to make the $\mathcal{S^{-}}$ more inappropriate, or $\mathcal{S^{+}}$ more appropriate. Similarly to Shao et al. salt-nlp2024privacylens, we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that would introduce reasoning shortcuts in the expansions. As we did for our label-dependent expansion of each field, for ConfAIde+ we again need to further integrate $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$. In total, we generate an additional $9\times 493=$ 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples for PrivacyLens+, and $9\times 270=$ 2,430 pairs for ConfAIde+. We provided detailed descriptions of our generation procedure in Appendix D.1.3 and  D.2.4.

### v3

@@ 1 Introduction
Personal AI agents are expected to interact with a user to complete tasks on their behalf. To effectively complete a broad variety of tasks in a personalized manner, such agents require access to significant amounts of personal and potentially sensitive user data. This implies that an agent must be able to safeguard this personal information while achieving high utility. To be truly useful, the agent must therefore be able to judge based on the context of the interaction when it is appropriate (as defined in the contextual integrity framework Nissenbaum (2004, 2009)) and consistent with a specific user’s expectations to share personal information in interactions with third-party tools and agents.
Prior research Mireshghallah et al. (2024); Shao et al. (2024) has investigated the ability of large language models (LLMs) to maintain the secrecy of sensitive information, with initial studies reporting limited success. More recent findings with advanced LLMs, however, indicate substantial improvements on certain datasets Sun et al. (2024). Nevertheless, these existing benchmarks focus on whether an LLM avoids disclosing sensitive information and do not equally address the crucial aspect of utility; that is, whether the model can contextually determine when sharing sensitive information is appropriate.
@@
First, we introduce a comprehensive dataset encompassing both positive (appropriate to share) and negative (inappropriate to share) examples, expanding upon existing datasets that predominantly feature negative instances. Our study then reveals an oversensitivity to prompt variations in LLM decisions—a signal that can indicate complexity in contextual integrity analysis Shvartzshnaider and Duddu (2024). We also find that classification errors frequently stem from underspecified context, leading the models to make implicit assumptions. In instances where model predictions diverge, the model’s internal reasoning is consistent with privacy reasoning we expect to see. However, it leads to a different outcome due to the implicit application of different, yet seemingly plausible, contextual interpretations. We term these scenarios “ambiguous contexts”: situations lacking the contextual details necessary to definitively determine the appropriateness of data sharing.
@@ 2 Background and Related Work
Reasoning and understanding. Nuanced privacy judgments benefit from an understanding of the mental states of other parties, through theory of mind. Such capabilities of LLMs have been studied both generally Kim et al. (2023); Sclar et al. (2023) and in relation to privacy Mireshghallah et al. (2024). Privacy assessments also rely on complex reasoning, which has been studied extensively through benchmarks in other domains. StrategyQA Geva et al. (2021) and MuSR Sprague et al. (2024) test implicit and multi-step reasoning. Logical reasoning errors have been explored in ZebraLogic Lin et al. (2025), the cause of reasoning errors has been investigated in REVEAL Jacovi et al. (2024), and recently the ability to ask clarifying questions in underspecified contexts has been studied in QuestBench Li et al. (2024) on logical puzzles, planning, and math problems. Our work complements benchmarks developed for general reasoning and increases the trustworthiness of these general-purpose models, through measurements of how advancements in general reasoning capabilities translate to improving privacy reasoning and assessments. It also extends the analysis of underspecified context to privacy reasoning.
@@
LLM privacy assessments. Prior work has studied privacy issues related to LLMs from multiple angles. One line of work aims to assess Carlini et al. (2021); Staab et al. (2024) and mitigate Abadi et al. (2016); Li et al. (2021); Yu et al. (2021) leakage of training data. As the capability of LLMs to consume larger contexts has improved Lewis et al. (2020), allowing them to perform agent tasks on user data Bagdasarian et al. (2024); Ngong et al. (2025); Zharmagambetov et al. (2025), the research area of inference-time privacy has emerged. Inference-time privacy assessments can generally be categorized as either studies of privacy leakage or privacy awareness Huang et al. (2025). Privacy leakage measurements, e.g. through red teaming, focus on capturing the refusal to answer Wang et al. (2023); Zhang et al. (2024); Sun et al. (2024), and do not assess utility in terms of sharing sensitive data in appropriate contexts. In contrast, privacy awareness assessments are based on probing-style prompts to test the capability of LLMs in various privacy settings, where the appropriateness of sharing data is context-dependent. Since the goal of our benchmark is to capture the impact of context underspecification on privacy assessments, we focus on privacy awareness.
@@
Privacy awareness and contextual integrity. Due to the contextual nature of appropriateness, many privacy awareness assessments of LLMs have been grounded in contextual integrity theory Nissenbaum (2004, 2009), which defines privacy as adherence to context-specific informational norms. Several studies focus on the alignment of models with norms and user preferences Fan et al. (2024); Li et al. (2025a); Shvartzshnaider et al. (2024). Multiple such studies focus on building benchmarks for various tasks. The work by Ghalebikesabi et al. Ghalebikesabi et al. (2025) develops a benchmark for form-filling. PrivaCI-Bench Li et al. (2025b) focuses on benchmarks for legal compliance. CI-Bench Cheng et al. (2024) builds a benchmark for assistant tasks. A concurrent study by Lan et al. Lan et al. (2025) focuses on improving privacy reasoning through reinforcement learning. These papers focus on judging the capability of LLMs in assessing information-sharing appropriateness by also measuring utility, unlike our study that focuses on evaluating the impact of underspecified context on privacy reasoning. Moreover, these benchmarks are either not available publicly, or are focused on specific domains such as law. The two publicly available benchmarks that we use in our work are PrivacyLens Shao et al. (2024) and ConfAIde Mireshghallah et al. (2024), which we describe in detail in Appendix D and expand upon in Section 4.
@@ 3 Problem Statement
We consider an agent, depicted in Figure 1, modeled by an LLM $\phi$, tasked with making a binary judgment about sharing a piece of information (e.g., a medical condition or an appointment time). Such information is served to the model through data structures called scenarios $\mathcal{S_{C}}$, parameterized by a set of contextual fields $\mathcal{C}$, often inspired by the contextual integrity theory Nissenbaum (2004, 2009). The task of the model $\phi$ is then to output $\phi(\mathcal{S_{C}})\in\{\texttt{yes},\texttt{no}\}$, where yes (respectively no) corresponds to the conclusion that it is appropriate (respectively not) to share the data in the scenario $\mathcal{S_{C}}$. See Figure 2 for an example.
@@ 4 Context Ambiguity Analysis
To better evaluate both the privacy and utility of LLMs in probing-style tasks, we use two existing privacy benchmarks: PrivacyLens Shao et al. (2024) and ConfAIde Mireshghallah et al. (2024). PrivacyLens contains 493 examples characterized by five contextual integrity fields: data type, data subject, data sender, data recipient and transmission principle. In ConfAIde, examples describe scenarios where an aware agent tries to disclose a sensitive personal detail about a subject agent to an oblivious agent for a reveal reason. Each of the 270 ConfAIde examples contains seven fields (the subject agent’s aware-agent-relationship and oblivious-agent-relationship, in addition to the five above), which are expanded into short descriptive stories on which privacy judgments are performed. Because the two datasets contain only inappropriate-to-share scenarios (denoted as $\mathcal{S^{-}}$) to measure privacy leakage, we augment them by incorporating instances where data is appropriate to share ($\mathcal{S^{+}}$). This allows us to perform comprehensive assessments of utility. For each $\mathcal{S^{-}}$, we generate a corresponding positive scenario $\mathcal{S^{+}}$ by modifying a subset of its fields. For PrivacyLens, we source these appropriate-to-share fields directly from the dataset, while for ConfAIde we generate them using an LLM. We validate the positive examples manually. The resulting augmented datasets, termed PrivacyLens+ and ConfAIde+, incorporate $\mathcal{S^{+}}$ alongside the original $\mathcal{S^{-}}$, and contain 986 and 540 examples respectively. A detailed description of the dataset generation procedure is found in Appendix D.
@@
Following existing benchmarks Mireshghallah et al. (2024), we start with a neutral prompt that asks “is it okay” to share the relevant data without providing more information or context. Next, we create a restrictive prompt variant by adding “Please answer the question considering privacy norms in the response” to the neutral prompt, akin to prior work Mireshghallah et al. (2024). Finally, we create a permissive prompt variant by adding “Please answer the question sharing as much as possible while respecting privacy.” The goal of these variants is to identify how performance changes across distinct input instructions, rather than to develop the best prompt variants for each setting. In addition to these variants that change privacy directives in the prompts, we also evaluate the model’s sensitivity to paraphrasing-style perturbations in Appendix B.2.4. The complete prompts are listed in Appendix E. For all experiments, we set temperature to 0 and the maximum output length to one token (the yes/no judgment).
@@
Table 1 highlights that model performance is highly sensitive to the prompt variant used across datasets. The neutral prompt generally yields high precision (85.6–92.2% for PrivacyLens+ and 98.6–100.0% for ConfAIde+) but lower recall (62.7–75.7% for PrivacyLens+ and 85.6–95.6% for ConfAIde+), indicating models prioritize privacy over utility, achieving high precision at the cost of low recall. The high precision also aligns with prior work finding that models more easily distinguish negative examples Sun et al. (2024).
@@ 5 Camber: A Context Disambiguation Framework
The first strategy serves as a baseline and adds arbitrary information to one of the data fields (e.g., clarifying the data subject of a request from “alumnus” to “an alumnus who had been recognized for contributions in environmental science”), helping us measure to what extent any additional privacy-agnostic information is useful. To ensure that the added information remains in agreement with the existing appropriateness label of the example, we prompt the model to generate only neutral information, in line with prior work Shao et al. (2024). In the second baseline strategy, we aim to synthetically add information that is more privacy-related and indicative of appropriateness to further resolve ambiguity (e.g., retrieving information that the student record was marked as “do not disclose”). To ensure that this added information remains in agreement with the appropriateness label, we include the label in the data generation prompt. We call the first strategy label-independent and the second label-dependent. Finally, our third strategy mimics an agent that reasons about which fields in the information request are ambiguous and seeks evidence to reduce this ambiguity (e.g., reasoning that the sender is not authorized to share alumni information externally). We call this the reasoning-guided expansion strategy (Section 5.2). Comparing results from the first two strategies reveals how important it is for agents to clarify privacy-specific aspects compared to asking general questions about the context. The reasoning-guided strategy further intends to shed light on what types of privacy relevant information would be most effective to acquire. To evaluate these strategies and their impact on improving privacy judgments, in Section 5.1 we describe the mechanisms enabling Camber to generalize across both the ConfAIde+ and PrivacyLens+ datasets. A detailed description of the Camber framework can be found in Appendix C.
@@ 5.1 Enabling context disambiguation across datasets
To investigate to what extent context that clarifies the appropriateness label is effective at aiding LLM judgment, we define a second baseline that adds contextual details to $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples according to the definition of the fields in each dataset. These label-dependent contexts are either inappropriate or appropriate, denoted as $\mathcal{N^{-}_{D}}$ and $\mathcal{N^{+}_{D}}$ respectively. Incorporating these contexts yields examples denoted as $\mathcal{S^{-}_{N^{-}_{D}}}$ and $\mathcal{S^{+}_{N^{+}_{D}}}$. For both datasets, we first ask the LLM to generate $\mathcal{N^{-}_{D}}$ (respectively, $\mathcal{N^{+}_{D}}$) for a specific field in $\mathcal{S^{-}}$ ($\mathcal{S^{+}}$) to make the example more inappropriate (respectively, appropriate). For ConfAIde+, we then additionally ask the LLM to integrate the generated context into the original example. It is important to ensure that the expansion does not result in reasoning shortcuts, wherein the added context directly reveals the appropriateness label. We achieve this by prompting the context-generating LLM to avoid directly expanding on the appropriateness label, and using heuristics that remove certain words indicative of the label, in line with prior work Shao et al. (2024). We repeat this procedure for all fields in each dataset, resulting in an additional 2,465 pairs of $\mathcal{S^{-}_{N^{-}_{D}}}$ and $\mathcal{S^{+}_{N^{+}_{D}}}$ examples for PrivacyLens+, and 1,890 pairs for ConfAIde+. The detailed generation procedure can be found in Appendix D.1.2 and D.2.3.
@@ 5.2 Reasoning-guided expansion
Building on extensive research into LLM reasoning Geva et al. (2021); Sprague et al. (2024); Lin et al. (2025); Jacovi et al. (2024) and factors influencing their outputs, we seek to understand appropriateness decisions of LLMs by prompting them to provide concise reasoning alongside each yes/no decision (our prompts can be found in Appendix E). Although elicited reasoning may not perfectly represent internal processes for models, it offers valuable insights into their decision-making Chen et al. (2025). This reasoning elicitation step yielded performance comparable to that of models not providing explicit reasoning (P=83.5%, R=74.4%, and $F_{1}$=78.7% for PrivacyLens+, and P=97.6%, R=75.9%, and $F_{1}$=85.4% for ConfAIde+).
To understand model reasoning and identify implicit assumptions, two authors analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; FPs were excluded due to insufficient samples). They examined the prompts and reasoning for those examples—blind to ground-truth labels and LLM judgments—to identify assumptions and develop the nine codes shown in Table 2. Subsequently, using this established codebook, the same two authors independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreement was resolved through discussion to achieve consensus Cascio et al. (2019). Table 2 presents the resulting code occurrences across datasets.
@@
To test if codes are more effective at reducing implicit reasoning assumptions, we prompt the LLM to generate label-dependent contexts based on each code (denoted as $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$) for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples respectively in both datasets. Our instruction specifically asks the LLM to first identify the most suitable field to expand based on the code and its definition, and then to expand the selected field following this code to make the $\mathcal{S^{-}}$ more inappropriate, or $\mathcal{S^{+}}$ more appropriate. Similarly to Shao et al. Shao et al. (2024), we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that would introduce reasoning shortcuts in the expansions. As we did for our label-dependent expansion of each field, for ConfAIde+ we again need to further integrate $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$. In total, we generate an additional $9\times 493=$ 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples for PrivacyLens+, and $9\times 270=$ 2,430 pairs for ConfAIde+. We provided detailed descriptions of our generation procedure in Appendix D.1.3 and  D.2.4.